# Tackling GenAI Copyright Issues: Originality Estimation and Generecization

### v3

@@ Abstract
The rapid progress of generative AI technology has sparked significant copyright concerns, leading to numerous lawsuits filed against AI developers. While various techniques for mitigating copyright issues have been studied, significant risks remain. Here, we propose a genericization method that modifies the outputs of a generative model to make them more generic and less likely to infringe copyright. To achieve this, we introduce a metric for quantifying the level of originality of data in a manner that is consistent with the legal framework. This metric can be practically estimated by drawing samples from a generative model, which is then used for the genericization process. As a practical implementation, we introduce PREGen, which combines our genericization method with an existing mitigation technique. Experiments demonstrate that our genericization method successfully modifies the output of a text-to-image generative model so that it produces more generic, copyright-compliant images. PREGen dramatically improves the performance of the existing method, reducing the likelihood of generating copyrighted characters by more than half when the names of copyrighted characters are used as the prompt. Additionally, although generative models have been found to generate copyrighted characters even when the names of characters are not directly mentioned in the prompt, PREGen almost completely suppresses the generation of copyrighted characters for such prompts.
@@ 1 Introduction
Generative models have exhibited performance that rivals human capabilities in creative tasks such as those involving image synthesis and language processing[rombach2022high, bubeck2023sparks]. However, this progress has also heightened concerns over copyright protection, leading to numerous lawsuits filed by creators against the AI developers[samuelson2023]. Copyright law aims to protect the rights of creators over their works, encouraging the creation of new works while maintaining the balance between the interests of creators and the public[litman1990]. Concerns about copyright issues with generative models arise primarily because these models are trained on massive datasets that often contain copyrighted works and are capable of producing outputs closely resembling their training data, potentially resulting in violation of exclusive rights of copyright owners[sag2023].
To mitigate copyright concerns, previous studies have explored methods for reducing the likelihood of producing outputs that resemble copyrighted training data by modifying the training or inference of generative models[vyas2023, chu2024, he2024fantasticcopyrightedbeastsnot]. Although some methods have been found effective in reducing the likelihood of generating copyrighted characters, a significant risk of generating them remains [he2024fantasticcopyrightedbeastsnot, lee2024talkin].
@@ 2 Quantifying originality
The capacity of generative models to be trained on a massive corpus of preexisting creations and learning the distribution of data can potentially be leveraged to measure the originality of works[tu2021, franceschelli2022, hacohen2024, hacohen2024similarities]. In particular, although the true $P(y|x)$ in Equation 1 is not generally available, generative models trained to learn the real-world data distribution can be used as an approximation in practice. For example, to quantify the originality of an image, a text-to-image generative model can be substituted as $P(y|x)$. In this case, $x$ might be thought of as the prompt that sets the distribution of the model.
@@ 4 Proof of concept
We use SDXL[podell2023sdxl], a text-to-image generative model, to obtain originality estimates (Equation 2) of two images of copyrighted characters, Mario from Super Mario Bros and Winnie the Pooh. We used five different prompts (Table 1) for each image as the conditioning $x$, designed to describe the characters at different levels of abstractness or specificity. The specific prompts were generated by ChatGPT and other prompts were formed by excerpting parts of these generated prompts. We also apply our genericization method to the outputs of the generative model produced with the specific prompts in Table 1.
@@
Abstract “a mustachioed character” “a bear”
Intermediate “a plump, mustachioed character dressed in brown shoes” “a plump, mustachioed character dressed in blue overalls” “a mustachioed character dressed in brown shoes and white gloves” “a chubby, golden-yellow bear” “a chubby bear with a red shirt” “a golden-yellow bear with a red shirt”
Specific “a plump, mustachioed character dressed in a red hat and shirt, blue overalls, brown shoes, and white gloves” “a chubby, golden-yellow bear with a red shirt”
@@
Images of copyrighted characters as well as the sample images produced by the generative model using each prompt are displayed in Figure 1. More specific prompts tend to produce images that share more visual elements with the copyrighted characters. See Appendix D for further details of the experimental settings.
@@
Our metric successfully assigns higher originality values when the generative model is given more abstract prompts and lower originality values with more specific prompts. This reflects the fact that typical generation with more specific prompts, which more closely describe the copyrighted images, is more likely to be highly similar to copyrighted images. In addition, originality estimates for images of copyrighted characters tend to be higher than typical output from the generative model, especially when more abstract prompts are used (Figure 2). However, our originality estimation method can fail when the distribution learned by the generative model is highly distorted (see Appendix A for a case of failure). PREGen discussed in Section 5 addresses this issue by introducing diversity in the prompts used for internally generated samples.
Figure 2: Originality estimates of copyrighted images and generated images. The panels on the left and right show the originality estimates of the images of Mario and Pooh, respectively. Estimated originality values of the copyrighted images (green bars with black whiskers representing standard deviation) are higher when more abstract prompts are used. Comparison with the mean and standard deviation of originality estimates of images produced by the generative model using each prompt (red dots and dotted lines) indicates that the originality of the copyrighted images tends to be substantially higher than typical outputs.
@@
Figure 3: Distribution of similarity before and after genericization. The panels on the left and right show the distributions of similarity for images generated by prompts associated with Mario and Pooh, respectively. Outputs concentrate more around images highly similar to generic images after genericization whereas outputs highly similar to the images of copyrighted characters are suppressed.
Samples $y_{i}$ with the lowest estimated originality among those selected as $y_{\textrm{generic}}$ as well as samples $y_{i}$ that are the most similar to the copyrighted images are displayed in Figure 4. All five samples (for each of the two specific prompt, respectively) that are the most similar (in other words, have the highest cosine similarity values) to the copyrighted images are not $y_{\textrm{generic}}$, meaning that they were produced internally by the model but not selected as the output. The five samples that are most similar to the copyrighted image of Mario among all samples produced using the prompt “a plump, mustachioed character dressed in a red hat and shirt, blue overalls, brown shoes, and white gloves” share some distinctive features (e.g. short stature, large round blue eyes, a round emblem on the front of the cap, a 3D graphic style rendering) with Mario, and those that are most similar to the copyrighted image of Pooh among all samples produced using the prompt “a chubby, golden-yellow bear with a red shirt” tend to share some distinctive features (e.g. human-like facial features with a less pronounced snout, a body shape and posture resembling a teddy bear) with Pooh. This qualitative observation supports our quantitative results, suggesting that our genericization method effectively suppressesu outputs similar to unique copyrighted images in a way that aligns well with human perception.
Figure 4: Generated images with low originality and high similarity to copyrighted characters. The top row shows five images each that have the lowest estimated originality among $y_{\textrm{generic}}$ generated using the prompt “a plump, mustachioed character dressed in a red hat and shirt, blue overalls, brown shoes, and white gloves” and those generated using the prompt “a chubby, golden-yellow bear with a red shirt,” respectively. The bottom row displays five images each that have the highest similarity to the images of Mario and Pooh, respectively, all of which were suppressed by applying the genericization method.
@@ 5 Practical implementation and perfomance
We propose an algorithm to mitigate copyright risks of generative models, PREGen (Prompt Rwriting-Enhanced Genericization), that combines our genericization method involving originality quantification with the standard prompt rewriting method 1.
@@
Instead of producing a single image using the clean prompt as in the standard prompt rewriting, PREGen first produces an ensemble of images and selects the most generic one as the output, leveraging our originality estimation method. The use of multiple different rewritten prompts in this process, which are similar to each other but not exactly the same, aims to overcome the issue of some generic keywords and phrases having strong association with copyrighted characters [he2024fantasticcopyrightedbeastsnot] (see also Appendix A, where we show that the phrase “Italian plumber” seems to be strongly associated with Mario). Note that rewritten prompts are produced in $b$ batches where $n$ prompts are generated within each batch (Algorithm 1). Within the same batch, all the previously generated prompts are fed into the language model $L$ to generate the next prompt, a set-up that was found to encourage diversity in the rewritten prompts (see the system prompt template in Appendix D).
@@
DETECT measures the similarity of generated images to copyrighted characters by using a detection system, typically a multimodal model such as GPT-4o, to determine whether a specific copyrighted character appears in the generated content. A lower DETECT score indicates that fewer copyrighted characters were generated, which helps in minimizing potential copyright infringement. CONS measures the consistency of the generated image with the user’s intent using VQAScore [lin2024evaluatingtexttovisualgenerationimagetotext]. It evaluates, in particular, whether the key characteristics (e.g. “cartoon mouse” for Mickey Mouse) requested by the user are present in the generated image. A higher CONS score indicates that the generated content aligns better with the user’s expectations, thereby improving user satisfaction. We use the same list of key characteristics used in the previous study [he2024fantasticcopyrightedbeastsnot]. Note that there is a natural trade-off between DETECT and CONS. In an extreme case, a generative model can refuse generation all together or generate a single harmless image regardless of what the prompt says, which would result in a perfect DETECT score. However, the user would likely not be satisfied with such a model. It is crucial to strike the right balance between DETECT and CONS for a generative model to be copyright safe and useful at the same time.
In addition to DETECT and CONS, to aid our analysis, we define p-CONS, a new metric that measures the alignment of the generation with the input prompt provided by the user. p-CONS can be considered as a more fine-grained measure of prompt consistency than the standard CONS. The exact definition of p-CONS and the details of the anlysis involving p-CONS can be found in Appendix B.
@@
“Batman” “An imposing figure draped in a dark, bat-inspired costume. His athletic build is accentuated by a black bodysuit and a cape, with a distinctive emblem of a bat across his chest. His face is obscured by a cowl, leaving only a grim mouth and sharp eyes visible. His utility belt holds an array of gadgets.”
“Captain America” “He is a muscular man, often seen in a patriotic uniform consisting of a star-spangled blue torso, red and white striped midsection, and a white star emblem on the chest. His helmet features an ’A’ symbol, and he wields an indestructible round shield of red, white, and blue.”
“Judy Hopps” “She is a petite anthropomorphic rabbit with a strong build and gray fur. Her eyes are large and purple, and her ears are long and pink inside. She typically wears a dark blue police uniform, along with a vest and knee pads. Her expressive face constantly reflects her determination and optimism.”
“Mario” “A short, plump man sporting a cap and overalls, both in a vibrant shade of blue. He wears a red shirt and gloves, and is always seen with a thick, bushy moustache. His rosy cheeks, bright blue eyes, and brown loafers complete his look.”
“Spider-man” “A superhero donning a full-body suit, predominantly red and blue with a web-like design. His mask covers his entire face, featuring white, reflective lenses rimmed with black. His emblem, a black spider, adorns his chest. He is known for his athletic build and acrobatic agility.”
@@
PREGen is implemented with $b=4$, $n=5$, and using CLIP as the distance metric $d$. We show in Appendix C the effect of changing $b$, ranging from $1$ to $4$. To avoid reintroducing copyright sensitive elements in the rewritten prompts $x_{ij}$, we use as $\tau$ a slightly modified version of $t$ by adding sentences that encourage diversity while retaining the core intent of the original prompt (see Appendix D).
@@ Results
PREGen tends to produce images that have some common characteristics as the images produced without any intervention, but without elemenents that are highly unique to copyrighted characters. Even when the standard method fails to exclude these elements, PREGen does succeed in doing so due to its advanced ability to genericize (Figure 5). In fact, analysis of p-CONS suggest that, although PREGen effectively reduces the risks of generating copyrighted characters and improves the likelihood of generating content that has the same key characteristics as the original prompt, the generated content may deviate more from the user’s original prompt than when the standard method is used (see Appendix B). This indicates that PREGen balances the trade-off between DETECT and CONS by generating outputs that have the same general characteristic as the original prompt, but with differences in the details.
@@
As we increase the number of batches generated internally, the performance of PREGen, especially the consistency with the intent of the prompt, tend to increase (see Appendix C). Although we report the results with $b=4$ in this section, PREGen achieves a performance comparable to or better than the standard method on DETECT and CONS metrics with $b=2$, which requires less computation.
@@ 6 Related works
Researchers have recently explored the application of computational methods to copyright legal issues. For example, a framework for testing substantial similarity based on Kolmogorov-Levin complexity has been proposed[scheffler2022]. The potential of machine learning to bring objectivity to the substantial similarity test has also been recognized in the legal literature[tu2021]. While a deep-learning-based method for measuring creativity has been suggested[franceschelli2022], such methods can only be effective for our purposes if they align with the framework of copyright law. More recently, the ability of generative models to capture the distribution of expressions in a massive dataset has been considered promising for measuring originality[hacohen2024, hacohen2024similarities]. In this context, a method involving text inversion and reconstructing the original image from the resulting text tokens has been discussed as a means to measure the originality of images[hacohen2024]. This paper contributes to the literature by introducing a general and interpretable mathematical formula for quantifying the level of originality, which is consistent with the legal framework and can be practically estimated using generative models.
@@
Previous studies have explored methods for reducing the likelihood of producing outputs that resemble copyrighted data by training an auxiliary generative model on non-copyrighted data and employing rejection sampling[vyas2023], modifying the training objectives of generative models[chu2024], and rewriting the prompts[he2024fantasticcopyrightedbeastsnot]. A genericization method was introduced in this paper, which significantly improves the performance of the existing prompt rewriting method. As an alternative approach for mitigating the risks of infringing copyright in training data, methods for computing fair shares of royalty and compensating the authors of training data have been proposed, where the shares are computed based on the contribution of copyrighted data to the likelihood of generation[deng2023, wang2024]. Such an approach might not be realistic when the generative model is trained on massive data scraped from the web, which is often the case.

### v4

@@ Abstract
The rapid progress of generative AI technology has sparked significant copyright concerns, leading to numerous lawsuits filed against AI developers. While various techniques for mitigating copyright issues have been studied, significant risks remain. Here, we propose a genericization method that modifies the outputs of a generative model to make them more generic and less likely to infringe copyright. To achieve this, we introduce a metric for quantifying the level of originality of data in a manner that is consistent with the legal framework. This metric can be practically estimated by drawing samples from a generative model, which is then used for the genericization process. As a practical implementation, we introduce PREGen, which combines our genericization method with an existing mitigation technique. Experiments demonstrate that our genericization method successfully modifies the output of a text-to-image generative model so that it produces more generic, copyright-compliant images. Compared to the existing method, PREGen reduces the likelihood of generating copyrighted characters by more than half when the names of copyrighted characters are used as the prompt, dramatically improving the performance. Additionally, while generative models can produce copyrighted characters even when their names are not directly mentioned in the prompt, PREGen almost entirely prevents the generation of such characters in these cases.
@@ 1 Introduction
Generative models have exhibited performance that rivals human capabilities in creative tasks such as those involving image synthesis and language processing [rombach2022high, bubeck2023sparks]. However, this progress has also raised concerns about copyright protection, leading to numerous lawsuits filed by creators against AI developers [samuelson2023]. Copyright law aims to protect the rights of creators over their works, encouraging the creation of new works while maintaining the balance between the interests of creators and the public [litman1990]. Concerns about copyright issues with generative models arise primarily because these models are trained on massive datasets that often contain copyrighted works and are capable of producing outputs closely resembling their training data, potentially resulting in violation of exclusive rights of copyright owners [sag2023].
To mitigate copyright concerns, previous studies have explored methods for reducing the likelihood of producing outputs that resemble copyrighted training data by modifying the training or inference of generative models [vyas2023, wang2024evaluating, he2024fantasticcopyrightedbeastsnot, ma2024datasetbenchmarkcopyrightinfringement]. Although some methods have been found effective in reducing the likelihood of generating copyrighted characters, a significant risk of generating them remains [he2024fantasticcopyrightedbeastsnot, lee2024talkin].
@@ 2 Quantifying originality
The capacity of generative models to be trained on a massive corpus of preexisting creations and learning the distribution of data can potentially be leveraged to measure the originality of works [tu2021, franceschelli2022, hacohen2024, hacohen2024similarities, haviv2024imageworththousandwords]. In particular, although the true $P(y|x)$ in Equation 1 is not generally available, generative models trained to learn the real-world data distribution can be used as an approximation in practice. For example, to quantify the originality of an image, a text-to-image generative model can be substituted as $P(y|x)$. In this case, $x$ might be thought of as the prompt that sets the distribution of the model.
@@ 4 Proof of concept
Evaluation protocal
We use SDXL [podell2023sdxl], a text-to-image generative model, to obtain originality estimates (Equation 2) of two images of copyrighted characters, Mario from Super Mario Bros and Winnie the Pooh. We used five different prompts (Table 1) for each image as the conditioning $x$, designed to describe the characters at different levels of abstractness or specificity. The specific prompts are designed to capture the essence of the visual aspects of the copyrighted characters, and other prompts were formed by excerpting parts of the specific prompts. We also apply our genericization method to the outputs of the generative model produced with the specific prompts in Table 1.
@@
Abstract “a mustachioed character" “a bear"
Intermediate “a plump, mustachioed character dressed in brown shoes" “a plump, mustachioed character dressed in blue overalls" “a mustachioed character dressed in brown shoes and white gloves" “a chubby, golden-yellow bear" “a chubby bear with a red shirt" “a golden-yellow bear with a red shirt"
Specific “a plump, mustachioed character dressed in a red hat and shirt, blue overalls, brown shoes, and white gloves" “a chubby, golden-yellow bear with a red shirt"
@@
Images of copyrighted characters as well as the sample images produced by the generative model using each prompt are displayed in Figure 1. More specific prompts tend to produce images that share more visual elements with the copyrighted characters. See Appendix F for further details of the experimental settings.
@@
Results
The results of originality estimation and genericization using CLIP are presented. The results with DINOv2 are similar and are shown in Appendix A.
@@ Results
Our metric successfully assigns higher originality values when the generative model is given more abstract prompts and lower originality values with more specific prompts. This reflects the fact that typical generation with more specific prompts, which more closely describe the copyrighted images, is more likely to be highly similar to copyrighted images. In addition, originality estimates for images of copyrighted characters tend to be higher than typical output from the generative model, especially when more abstract prompts are used (Figure 2). However, our originality estimation method can fail when the distribution learned by the generative model is highly distorted (see Appendix C for a case of failure). PREGen discussed in Section 5 addresses this issue by introducing diversity in the prompts used for internally generated samples.
Figure 2: Originality estimates of copyrighted images and generated images (CLIP). The panels on the left and right show the originality estimates of the images of Mario and Pooh, respectively. Estimated originality values of the copyrighted images (green bars with black whiskers representing standard deviation) are higher when more abstract prompts are used. Comparison with the mean and standard deviation of originality estimates of images produced by the generative model using each prompt (red dots and dotted lines) indicates that the originality of the copyrighted images tends to be substantially higher than typical outputs.
@@
Figure 3: Distribution of similarity before and after genericization (CLIP). The panels on the left and right show the distributions of similarity for images generated by prompts associated with Mario and Pooh, respectively. Outputs concentrate more around images highly similar to generic images after genericization whereas outputs highly similar to the images of copyrighted characters are suppressed.
Samples $y_{i}$ with the lowest estimated originality among those selected as $y_{\textrm{generic}}$ as well as samples $y_{i}$ that are the most similar to the copyrighted images are shown in Appendix B. Visual elements that are highly unique to copyrighted characters tend to be absent from genericized sample, which supports the quantitative results that our genericization method effectively makes the output more generic.
@@ 5 Practical implementation and perfomance
We propose an algorithm to mitigate copyright risks of generative models (Algorithm 1), PREGen (Prompt Rwriting-Enhanced Genericization), [^2] The code is available at: https://github.com/hirochok/PREGen. that combines our genericization method involving originality quantification with the standard prompt rewriting method.
@@
Instead of producing a single image using the clean prompt as in the standard prompt rewriting, PREGen first produces an ensemble of images and selects the most generic one as the output, leveraging our originality estimation method. The use of multiple different rewritten prompts in this process, which are similar to each other but not exactly the same, aims to overcome the issue of some generic keywords and phrases having strong association with copyrighted characters [he2024fantasticcopyrightedbeastsnot] (see also Appendix C, where we show that the phrase “Italian plumber" seems to be strongly associated with Mario). Note that rewritten prompts are produced in $b$ batches where $n$ prompts are generated within each batch (Algorithm 1). Within the same batch, all the previously generated prompts are fed into the language model $L$ to generate the next prompt, a set-up that was found to encourage diversity in the rewritten prompts (see the system prompt template in Appendix F).
@@
DETECT measures the similarity of generated images to copyrighted characters by using a detection system, typically a multimodal model such as GPT-4o, to determine whether a specific copyrighted character appears in the generated content. A lower DETECT score indicates that fewer copyrighted characters were generated, which helps in minimizing potential copyright infringement. CONS measures the consistency of the generated image with the user’s intent using VQAScore [lin2024evaluatingtexttovisualgenerationimagetotext]. It evaluates, in particular, whether the key characteristics (e.g. “cartoon mouse" for Mickey Mouse) requested by the user are present in the generated image. A higher CONS score indicates that the generated content aligns better with the user’s expectations, thereby improving user satisfaction. We use the same list of key characteristics used in the previous study [he2024fantasticcopyrightedbeastsnot]. Note that there is a natural trade-off between DETECT and CONS. In an extreme case, a generative model can refuse generation all together or generate a single harmless image regardless of what the prompt says, which would result in a perfect DETECT score. However, the user would likely not be satisfied with such a model. It is crucial to strike the right balance between DETECT and CONS for a generative model to be copyright safe and useful at the same time.
In addition to DETECT and CONS, to aid our analysis, we define p-CONS, a new metric that measures the alignment of the generation with the input prompt provided by the user. p-CONS can be considered as a more fine-grained measure of prompt consistency than the standard CONS. The exact definition of p-CONS and the details of the anlysis involving p-CONS can be found in Appendix D.
@@
“Batman" “An imposing figure draped in a dark, bat-inspired costume. His athletic build is accentuated by a black bodysuit and a cape, with a distinctive emblem of a bat across his chest. His face is obscured by a cowl, leaving only a grim mouth and sharp eyes visible. His utility belt holds an array of gadgets."
“Captain America" “He is a muscular man, often seen in a patriotic uniform consisting of a star-spangled blue torso, red and white striped midsection, and a white star emblem on the chest. His helmet features an ’A’ symbol, and he wields an indestructible round shield of red, white, and blue."
“Judy Hopps" “She is a petite anthropomorphic rabbit with a strong build and gray fur. Her eyes are large and purple, and her ears are long and pink inside. She typically wears a dark blue police uniform, along with a vest and knee pads. Her expressive face constantly reflects her determination and optimism."
“Mario" “A short, plump man sporting a cap and overalls, both in a vibrant shade of blue. He wears a red shirt and gloves, and is always seen with a thick, bushy moustache. His rosy cheeks, bright blue eyes, and brown loafers complete his look."
“Spider-man" “A superhero donning a full-body suit, predominantly red and blue with a web-like design. His mask covers his entire face, featuring white, reflective lenses rimmed with black. His emblem, a black spider, adorns his chest. He is known for his athletic build and acrobatic agility."
@@
PREGen is implemented with $b=4$, $n=5$, and using CLIP as the distance metric $d$. We show in Appendix E the effect of changing $b$, ranging from $1$ to $4$. To avoid reintroducing copyright sensitive elements in the rewritten prompts $x_{ij}$, we use as $\tau$ a slightly modified version of $t$ by adding sentences that encourage diversity while retaining the core intent of the original prompt (see Appendix F).
@@ Results
PREGen tends to produce images that have some common characteristics as the images produced without any intervention, but without elemenents that are highly unique to copyrighted characters. Even when the standard method fails to exclude these elements, PREGen does succeed in doing so due to its advanced ability to genericize (Figure 4). In fact, analysis of p-CONS suggest that, although PREGen effectively reduces the risks of generating copyrighted characters and improves the likelihood of generating content that has the same key characteristics as the original prompt, the generated content may deviate more from the user’s original prompt than when the standard method is used (see Appendix D). This indicates that PREGen balances the trade-off between DETECT and CONS by generating outputs that have the same general characteristic as the original prompt, but with differences in the details.
@@
As we increase the number of batches generated internally, the performance of PREGen, especially the consistency with the intent of the prompt, tend to increase (see Appendix E). Although we report the results with $b=4$ in this section, PREGen achieves a performance comparable to or better than the standard method on DETECT and CONS metrics with $b=2$, which requires less computation.
@@ 6 Related works
Generative AI and copyright law
Recently, a substantial body of literature has emerged exploring the intersection of copyright law and generative AI [sag2023, lemley2024, abbott2023disrupting, lee2024talkin, henderson2023, cooper2024filescomputercopyrightmemorization, samuelson2024, yang2024generativeaicopyrightdynamic, desai2024copyrightcomputersciencelaw, pasquale2024]. While we have addressed certain aspects of copyright law as they pertain to our work, we refer readers to these sources for a more comprehensive discussion of the various issues involved.
@@
Researchers have recently explored the application of computational methods to copyright legal issues. For example, a framework for testing substantial similarity based on Kolmogorov-Levin complexity has been proposed [scheffler2022]. The potential of machine learning to bring objectivity to the substantial similarity test has also been recognized in the legal literature [tu2021]. While a deep-learning-based method for measuring creativity has been suggested [franceschelli2022], such methods can only be effective for our purposes if they align with the framework of copyright law. More recently, the ability of generative models to capture the distribution of expressions in a massive dataset has been considered promising for measuring originality [hacohen2024, hacohen2024similarities, haviv2024imageworththousandwords]. In this context, a method involving text inversion and reconstructing the original image from the resulting text tokens has been discussed as a means to measure the originality of images [hacohen2024, haviv2024imageworththousandwords]. This paper contributes to the literature by introducing a general and interpretable mathematical formula for quantifying the level of originality, which is consistent with the legal framework and can be practically estimated using generative models.
@@
Previous studies have explored methods for reducing the likelihood of producing outputs that resemble copyrighted data by ensuring that the output distribution of the model is close to a model trained without access to copyrighted content [vyas2023], refusing the outputs of detected copyrighted content and guiding the generation away from such content [wang2024evaluating], removing specific content from generative models via application of various unlearning methods [ma2024datasetbenchmarkcopyrightinfringement], and rewriting the prompts [he2024fantasticcopyrightedbeastsnot]. A method for genericizing outputs by quantifying originality was introduced in this paper, which significantly improves the performance of the existing prompt rewriting method. While we have focused on image generation, methods for rejecting or modifying copyright-sensitive outputs of language models have also been proposed and evaluated in the literature [chu2024, min2024silolanguagemodelsisolating, wei2024evaluatingcopyrighttakedownmethods, dou2024avoidingcopyrightinfringementmachine, abad2024strongcopyrightprotectionlanguage, chen2024copybenchmeasuringliteralnonliteral]. As an alternative approach for mitigating the risks of infringing copyright in training data, methods for computing fair shares of royalty and compensating the creators of training data have been proposed, where the shares are computed based on the contribution of copyrighted data to the likelihood of generation [deng2023, wang2024].